OpenStudy (b.r):

which property is illustrated by the following statement? 6 + (2x + 2x) = (6 + 2x) + 2z

OpenStudy (b.r):

@Directrix

Directrix (directrix):

Any idea what it might be?

Directrix (directrix):

Is one of the x terms on the left supposed to be a z? 6 + (2x + 2x) = (6 + 2x) + 2z

OpenStudy (anonymous):

commulative I believe. However you spell it.

OpenStudy (b.r):

I thought communicative property.

Directrix (directrix):

6 + (2x + 2z) = (6 + 2x) + 2z Do you see that you posted the problem incorrectly? One side has a z and the other does not.

OpenStudy (anonymous):

If you meant this problem ^, you are correct I believe.

Directrix (directrix):

This is how it was posted. See attachment.

Directrix (directrix):

The commutative property is the order property. The associative property is the grouping property.

Directrix (directrix):

6 + (2x + 2z) = (6 + 2x) + 2z (6 + 2x ) + z = (6 + 2x) + 2z

OpenStudy (b.r):

Yeah. I meant the x is a z.

Directrix (directrix):

The order of the terms on both sides is 6, 2x, 2z So, the answer is not commutative.

Directrix (directrix):

If this is the problem: 6 + (2x + 2z) = (6 + 2x) + 2z then the answer is the grouping property. What is its name?

OpenStudy (b.r):

Accoustative property of addition?

OpenStudy (b.r):

Something like that.

Directrix (directrix):

The associative property is the grouping property. That is correct. Associative property. @B.R

OpenStudy (b.r):

I have another question.

Directrix (directrix):

What is it?

OpenStudy (b.r):

What is the simplified form of this expression?

OpenStudy (b.r):

@Directrix

Directrix (directrix):

That means: 3/5 * 3/5 * 3/5 = ?

Directrix (directrix):

What is 3 * 3*3 ? That is the numerator. What is 5 * 5 * 5 ? That is the denominator.

Directrix (directrix):

@B.R

OpenStudy (b.r):

@Directrix God- I am stuck on this one.

Directrix (directrix):

What is 3 times 3 times 3? Post what you get.

Directrix (directrix):

You can paste this into the Google search box and the Google calculator will pop up with the product for you. 3*3*3=

Directrix (directrix):

After that, do this on Google. 5 * 5 * 5 =

OpenStudy (b.r):

27

OpenStudy (b.r):

125